OpenStudy (anonymous):

Find the center and the radius of each circle. Graph each circle. Describe the translation from center (0, 0). (x + 7)^2 + (y + 3)^2 = 49 I found the center just need help on radius

OpenStudy (anonymous):

Circles take the form \[\Large (x-h)^2 + (y-k)^2= r^2\] Where the center is (h,k) and the radius is r ^_^ Follow the pattern :D

OpenStudy (anonymous):

(x-(-7)^2+ (y-(-3)^2

OpenStudy (anonymous):

That's good, so you know how to get the center, no problem ^^ There's just one little thing, and that's the radius. In your particular equation, which value corresponds to \(\large r^2\) ?

OpenStudy (anonymous):

I have no idea..

OpenStudy (anonymous):

Let me just lay it down for you, for your convenience ^^ \[\Large (x - \color{green}h)^2 + (y-\color{red}k)^2 = \color{blue}r^2\]\[\Large (x\color{green}{+7})^2 + (y\color{red}{+3})^2 = \color{blue}{49}\] better? :D

OpenStudy (anonymous):

So... *now* do you see which value corresponds to \(\large r^2\) ? It's staring you right in the face.... ;)

OpenStudy (anonymous):

49?

OpenStudy (anonymous):

Yup. And BE CAREFUL, that's not the radius, that's the SQUARE of the radius. Given that, what's the radius r, finally? :D

OpenStudy (anonymous):

7? 24.5?

OpenStudy (anonymous):

Which does your heart tell you? ;) Remember... 49 is the square of the radius... the squaaare of the raaaadius (this works much better in a ghostly spooky voice)

OpenStudy (anonymous):

7? Hahaha I dont know. I am drawing a blank.!

OpenStudy (anonymous):

Trust your heart, Texas Girl (please feel free to tell me how you prefer to be called lol) It can't be 24.5, that's way off, if you square 24.5, you get 600.25... which is so not 49 :D \(\large 7^2\) on the other hand,is exactly 49. So, are your heart and mind at ease now? ^^

OpenStudy (anonymous):

Yes! Thank you so much!!

OpenStudy (anonymous):

Good. My work here is done... [Kurt walks off into the sunset]

OpenStudy (anonymous):

I assume, now that you know the radius and the center, graphing this blasted circle would be a no-brainer? :D

OpenStudy (anonymous):

Yes. Thanks a lot!

OpenStudy (anonymous):

Okay. Have fun :D